OpenStudy (anonymous):

I have a question! (WILL GIVE MEDALS) How do you find the half of 6 1/6

OpenStudy (anonymous):

there that will help

OpenStudy (anonymous):

6 1/6 = 6 + 1/6 = 37/6 Could you find the half of 37/6 ?

OpenStudy (anonymous):

18/3 i think

OpenStudy (anonymous):

OH i made a mistake thnk you all for your help ;)

OpenStudy (anonymous):

\[\frac{ 37 }{ 6 } \times \frac{ 1 }{ 2 } = ?\]

OpenStudy (hoblos):

a(b+c) = ab+ac \[\frac{ 1 }{ 2 }(6+\frac{ 1 }{ 6 }) = \frac{ 1 }{ 2 }6 + \frac{ 1 }{ 2 }\frac{ 1 }{ 6 }\]

OpenStudy (anonymous):

37 times 1 = 37 6 times 2 = 12 37/12

OpenStudy (anonymous):

Thank you Hoblos that realy helped.

OpenStudy (hoblos):

so you got the answer ?

OpenStudy (anonymous):

GIve me one sec.

OpenStudy (anonymous):

1/2 + 1/2 + 1/2 + 1/2 + 1/2 + 1/2 + 1/2 + 1/2 + 1/2 + 1/2 + 1/2 + 1/2 + 1/2 = 6 1/2

OpenStudy (anonymous):

it goes 13 times